OpenStudy (anonymous):

can someone PLEASE explain to me what scientific notations are and how to use them?!?!

OpenStudy (anonymous):

scientific notation example would be if you have 10000 in scientific notation it would be 1x10^4 <-- the 4 as how many 0's after 1

OpenStudy (anonymous):

so if you have 523 and you want to write it in scientific notation then it would be 5.23x10^2

OpenStudy (ranga):

In scientific notation the numbers are written in the form: M x 10^N Where M is a a number in the range: 1 <= M < 10. M can be an integer or a decimal. N is an integer. Regarding usage, let us say you are asked to multiply 5000 and 80000. First put 5000 in scientific notation: 5000 = 5 x 10^3 Put 80000 in scientific notation: 80000 = 8 x 10^4 5000 x 80000 = 5 x 10^3 * 8 x 10^4 = 40 x 10^(3 + 4) = 40 x 10^7 = 4 x 10^8